Please answer this equation:
Find the equation of the line that goes through the point (3, 2) and has a slope of -3 over 2.
Thanks (:

Answer by sudhanshu_kmr(1152)


y-2 = -3/2 * (x-3)
=> 2y -4 = -3x + 9
=> 2y = -3x + 13
=> y = -3x/2 + 13/2